40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 8 20:11:54 2012, Jenne, Update, Alignment, AS steering moved out of POY's way, 2 green beams onto PSL table AS_8Nov2012.pngREFL_8Nov2012.png
    Reply  Thu Nov 8 20:54:08 2012, Manasa, Ayaka, Update, Alignment, Reconfirming on IPPOS, IPANG and oplevs centering 
    Reply  Fri Nov 9 08:11:57 2012, Steve, Update, Alignment, nothing moved - all good opsingGOOD.pngop8h.png8amF.pngkeepfailing.png
       Reply  Fri Nov 9 11:38:38 2012, jamie, Update, General, we're closing up 
          Reply  Fri Nov 9 17:15:05 2012, Manasa, Steve, Ayaka, Update, General, We're closed! Pumping down monday morning 
Message ID: 7694     Entry time: Fri Nov 9 17:15:05 2012     In reply to: 7693
Author: Manasa, Steve, Ayaka 
Type: Update 
Category: General 
Subject: We're closed! Pumping down monday morning 

Quote:

After a brief look this morning, I called it and declared that we were ok to close up.  The access connector is almost all buttoned up, and both ETM doors are on.

Basically nothing moved since last night, which is good.  Jenne and I were a little bit worried about how the input pointing might have been effected by our moving of the green periscope in the MC chamber.

First thing this morning I went into the BS chamber to check out the alignment situation there.  I put the targets on the PRM and BS cages.  We were basically clear through the PRM aperture, and in retro-reflection.

The BS was not quite so clear.  There is a little bit of clipping through the exit aperture on the X arm side.  However, it didn't seem to me like it was enough to warrant retouching all the input alignment again, as that would have set us back another couple of days at least.

Both arm green beams are cleaning coming out, and are nicely overlapping with the IR beams at the BS (we even have a clean ~04 mode from the Y arm).  The AS and REFL spots look good.  IPANG and IPPOS are centered and haven't moved much since last night.  We're ready to go.

The rest of the vertex doors will go on after lunch.

Jamie and Steve got the ETM doors on this morning.

We got the other heavy doors including the ITMs, BS and the access connector in place.

If nobody raises any concerns in reply to this elog, Steve will assume it as a green signal and will start pumping down first thing Monday morning after the final check on the access connector bellow screws.

 

Steve! 

Ayaka and I got the ITMY and BS door closed at 45foot pounds just now. 

ELOG V3.1.3-